OpenStudy (anonymous):

if the average of 14 consecutive integers is 20.5, what is the average of the first seven integers?

OpenStudy (anonymous):

Let \(n\) be an integer, then \[\large\frac{\displaystyle\sum_{k=0}^{13}(n+k)}{14}=20.5\] You should be able to solve for \(n\) at this point using some known summation formulas: \[\large\sum_{k=1}^r1=r\\ \large\sum_{k=1}^r k=\frac{r(r+1)}{2}\] You have \[\large\begin{align*}\sum_{k=0}^{13}(n+k)&=n+\sum_{k=1}^{13}(n+k)\\ &=n+n\sum_{k=1}^{13}1+\sum_{k=1}^{13}k\\ &=n+13n+\frac{13(13+1)}{2}\\ &=14n+91\end{align*}\] so the first equation has become \[\large\frac{14n+91}{14}=20.5~~\iff~~n=14\] The average of the first seven integers would then be \[\frac{14+15+16+17+18+19+20}{7}=\cdots\]